Suppose a Monopolist faces the following Total Cost and Demand functions:


TC = 100 +2Q


P = 25 - 0.25 What is the firm’s profit-maximizing position?



1
Expert's answer
2022-03-13T18:53:35-0400

For profit maximization, MR= MC

TC = 100 +2Q

MC= 2Q

P = 25 - 0.25Q

TR= P. Q

= Q(25-0.25Q)

= 25Q+ 0.25"Q^2"

MR= 25- 0.5Q

"\therefore 25-0.5Q= 2Q"

2.5Q= 25

Q= 10

P= 25- 0.25(10)

= 25- 2.5= 22.5